the square of a certain number minus nine times the number is 36. find the number.
Answer by stanbon(75887)

Let the number be "x".
Then x^2-9x=36
x^2-9x-36=0
(x-12)(x+3)=0
x=12 or x=-3
Cheers,
Stan H.